# RLCAD: Reinforcement Learning Training Gym for Revolution Involved CAD Command Sequence Generation

### v2

@@ 5.4 Policy Network
Our policy network consists of feature extraction modules and a Critic-Actor network. As shown in Figure  2, the network processes three inputs: B-Reps of the target and current models, and historical action sequences. First, we extract B-Rep modeling of the target and current models, encoding each B-Rep graph through an independent UV-Net to produce 256-dimensional embeddings $g_{t}$ (target) and $g_{c}$ (current). These embeddings are stacked into $G_{stack}\in\mathbb{R}^{2\times 256}$, which undergoes 8-head self-attention alignment, where $G_{stack}$ serves as both Query and Key/Value inputs. This generates the refined tensor $\tilde{G}_{stack}$. We then split $\tilde{G}_{stack}$ along the stacking dimension to obtain enhanced features $\tilde{g}_{t}$ and $\tilde{g}_{c}$, finally concatenating them into a fused geometric representation $G_{cat}\in\mathbb{R}^{512}$.
For action sequence processing, historical actions $(a_{0},a_{1},\dots,a_{t})$ are first embedded into 256-dimensional vectors with sinusoidal positional encodings, then fed into an 8-head Gated Transformer-XL (GTrXL) [43] module composed of $L$ stacked blocks. Each GTrXL block stabilizes the modeling of long-range dependencies through a gating mechanism. Finally, we extract the temporal features $h_{\text{last\_valid}}^{(L)}$ from the hidden state of the last valid action in the $L$-th layer. We empirically choose the number of stacked layers $L=3$.
@@ 5.5 Reward Shaping
However, observations reveal that relying solely on IoU led to suboptimal policies. The agent prioritized maximizing coarse volumetric overlap while neglecting fine geometric details. To address this limitation, we introduce three complementary rewards: Minimum Matching Distance (MMD) [44], Normal Consistency (NC) [45] and Neural Reward (NR) which explicitly enforce local geometric fidelity, surface quality and high-dimensional feature similarity.
Minimum Matching Distance (MMD) quantifies the average distance between the generated model and its closest-matching reference shape. It leverages Chamfer Distance (CD) and Earth Mover’s Distance (EMD) [44] to measure bidirectional geometric discrepancies. For two point clouds $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ and $\mathcal{Y}=\{y_{j}\}_{j=1}^{M}$sampled from the surface:
@@
[…]
@@ 6 Experiments and Comparison
We construct a CAD dataset containing extrusion and revolution operations for training and evaluation, which includes $20k$ CAD models from the ABC dataset  [27] after online parsing and deduplication. The dataset is stratified by the number of faces to measure model complexity, comprising $14k$ simple models (with fewer than 10 faces), $4k$ medium models (with 10 to 20 faces), and $2k$ complex models (with more than 20 faces). Using stratified sampling, we divide the dataset into a training set of $19k$ models and a testing set of $1k$ models, where complex models with over 20 faces account for 39.8% of the test dataset.
To further enhance the pre-training performance, we employ the DeepCAD dataset ($170k$ samples) in conjunction with our training dataset ($19k$ samples), forming a comprehensive pre-training corpus of $189k$ CAD models. We analyze the geometric complexity of the pre-training dataset in terms of face count per model: 33% contain fewer than 10 faces, 32% have 10–20, 12% have 20–30, 6% have 30–40, and 17% contain more than 40 faces.
@@ 6.1 Evaluation Metrics
To assess the quality of the generated CAD models, we first convert the B-Rep models into point clouds. Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models. We employ IoU, COV, MMD-CD, JSD, and NC as metrics to evaluate the quality of our generated CAD models [44]. IoU, MMD-CD, and NC have been introduced in Section 5.5.
@@ 6.2 Training Efficiency Comparison
Fusion 360 provides a server-side plugin called Fusion 360 Gym [11], which sequentially receives CAD modeling commands sent from external programs and returns the results to those programs via Hypertext Transfer Protocol (HTTP). Based on this plugin, we connect our reinforcement learning algorithm with Fusion 360 to enable closed-loop training.
We compared our novel training gym with Fusion 360 Gym over 10,000 training steps. Using parallel computation (1, 8, and 16 environments), we evaluated both memory consumption and execution time per step. Table 1 shows that Fusion 360 Gym consumes substantially more memory. The frequent use of the revert operation in Fusion 360 introduces a risk of memory overflow. In contrast, our gym employs a stable mark-and-revert mechanism that maintains steady memory usage without significant fluctuations. In addition, by replacing Fusion 360’s network communication mechanism with multi-process parallelism and shared memory architecture, we achieve a significant reduction in data transmission latency between the reinforcement learning algorithm and the gym. We get up to $39X$ speedup compared with Fusion 360 Gym, as shown in Table 2.
@@
Table 2: Comparison of execution time per step (unit: ms). It shows that our method get up to $39X$ speedup compared with Fusion 360 Gym.
@@ 6.3 B-Rep Modeling Comparison
    Graph Transformer Network (GTN) [46]: Utilize multi-head self-attention with edge-aware positional encoding, effectively capturing long-range dependencies and global topological constraints.
@@ 6.5 Reward Ablation Study
Table 5: Reward function ablation study. It shows that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, while metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
@@
Leveraging the ablated dataset curated in Section 6.3, we perform ablation studies to systematically analyze the impact of different reward function combinations on experimental results. Table 5 shows the outcomes of these studies. Reward functions such as IoU, MMD, and NC are used to measure geometric similarity. NR is used to capture the cosine similarity of high-dimensional B-Rep features. The experimental results show that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, meanwhile, metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
@@ 6.6 Result Comparison of Reconstruction
Table 6: Quantitative results on the B-Rep-based reconstruction task. Our method significantly outperforms Fusion 360 Gallery across all metrics.
@@
Table 7: Quantitative comparison with cadrille and CAD-Recode on the reconstruction task. Our method significantly outperforms these two methods across all metrics.
@@
    Fusion 360 Gallery [11] extracts continuous geometric parameters from B-Rep models and uses imitation learning to generate command sequences, followed by a confidence-based local search for optimization.
@@
    CAD-Recode [47] translates a point cloud into Python code by leveraging a relatively small language model as a decoder and combined with a lightweight point cloud projector. The method is trained on a self-constructed, million-scale training dataset composed exclusively of sketch-extrusion operations.
@@
    cadrille [48] proposes a multi-modal CAD reconstruction model that can simultaneously process three input modalities, including point clouds, images, and text. The method adopts a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning fine-tuning using online feedback.
Currently, most supervised learning methods are trained on datasets that include only sketch-extrusion operations. To overcome this limitation, we extend the output modules of Fusion 360 Gallery models to support revolve operations. And we retrain the model on our dataset. As illustrated in Figure 6, our approach outperforms Fusion 360 Gallery in completeness and detail preservation. Quantitative results in Table 6 demonstrate substantial improvements across all evaluation metrics.
@@ 6.7 Failure Case
We categorize the cases that our method currently cannot handle into three types. The first type arises from complex model with many-faces. Figure 8a exhibits the generation result of a complex gear model, in which some details is missing. Figure 9 depicts the quantitative relationship between generated model complexity (measured by face count) and various evaluation metrics. For simple models with a low face count, IoU and NC remain relatively high, but gradually decline as complexity increases. COV performs well on models with a low number of faces but declines significantly once the face count exceeds approximately 35–40. JSD, which is initially low, rises rapidly as the face count increases, especially beyond 40 faces. Overall, while the global reconstruction quality of our model remains acceptable beyond 35–40 faces, the fidelity of fine-grained geometric details deteriorates noticeably.
@@ 7.2 Future Work
Through extensive experiments, it has been found that our method still has shortcomings in detailed feature learning of complex models and supported operation amount. In the next phase, We will focus on developing detail-preserving feature extraction networks that support a broader range of input modalities, and expanding the gym to support more commonly used modeling operations.

### v3

@@ 5.4 Policy Network
Our policy network consists of feature extraction modules and a Critic-Actor network. As shown in Figure 2, the network processes three inputs: B-Reps of the target and current models, and historical action sequences. First, we extract B-Rep modeling of the target and current models, encoding each B-Rep graph through an independent UV-Net to produce 256-dimensional embeddings $g_{t}$ (target) and $g_{c}$ (current). These embeddings are stacked into $G_{stack}\in\mathbb{R}^{2\times 256}$, which undergoes 8-head self-attention alignment, where $G_{stack}$ serves as both Query and Key/Value inputs. This generates the refined tensor $\tilde{G}_{stack}$. We then split $\tilde{G}_{stack}$ along the stacking dimension to obtain enhanced features $\tilde{g}_{t}$ and $\tilde{g}_{c}$, finally concatenating them into a fused geometric representation $G_{cat}\in\mathbb{R}^{512}$.
For action sequence processing, historical actions $(a_{0},a_{1},\dots,a_{t})$ are first embedded into 256-dimensional vectors with sinusoidal positional encodings, then fed into an 8-head Gated Transformer-XL (GTrXL) [parisotto2020stabilizing] module composed of $L$ stacked blocks. Each GTrXL block stabilizes the modeling of long-range dependencies through a gating mechanism. Finally, we extract the temporal features $h_{\text{last\_valid}}^{(L)}$ from the hidden state of the last valid action in the $L$-th layer. We empirically choose the number of stacked layers $L=3$.
@@ 5.5 Reward Shaping
However, observations reveal that relying solely on IoU led to suboptimal policies. The agent prioritized maximizing coarse volumetric overlap while neglecting fine geometric details. To address this limitation, we introduce three complementary rewards: Minimum Matching Distance (MMD) [achlioptas2018learning], Normal Consistency (NC) [yu2022monosdf], and Neural Reward (NR), which explicitly enforce local geometric fidelity, surface quality, and high-dimensional feature similarity.
Minimum Matching Distance (MMD) quantifies the average distance between the generated model and its closest-matching reference shape. It leverages Chamfer Distance (CD) and Earth Mover’s Distance (EMD) [achlioptas2018learning] to measure bidirectional geometric discrepancies. For two point clouds $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ and $\mathcal{Y}=\{y_{j}\}_{j=1}^{M}$sampled from the surface:
@@
$\text{MMD-CD}(\mathcal{G},\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{\mathcal{Y}\in\mathcal{S}}\min_{\mathcal{X}\in\mathcal{G}}d_{CD}(\mathcal{X},\mathcal{Y}),$
$\text{MMD-EMD}(\mathcal{G},\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{\mathcal{X}\in\mathcal{G}}\min_{\mathcal{Y}\in\mathcal{S}}d_{EMD}(\mathcal{X},\mathcal{Y}).$
@@ 6 Experiments and Comparison
We construct our datasets from two primary sources: 170kk models from the DeepCAD dataset and a custom-curated corpus of 20kk models drawn from the ABC dataset [koch2019abc]. To construct this custom corpus, we first conducted online parsing and deduplication, retaining only models involving extrusion and revolution operations. From this filtered pool, we then sampled 20kk models using random stratified sampling based on geometric complexity categories: simple (fewer than 10 faces), medium (10–20 faces), and complex (more than 20 faces).
From the custom corpus of 20kk models, we allocated a 1kk-model test dataset for final evaluation using a random stratified sampling strategy, where 39.8% of models contain more than 20 faces. The remaining 19kk models were combined with the 170kk DeepCAD models to form a 189kk-model pre-training dataset. This pre-training dataset covers a wide range of geometric complexities: $33\%$ of models have fewer than 10 faces, $32\%$ have 10–20 faces, and $35\%$ have more than 20 faces.
@@ 6.1 Evaluation Metrics
To assess the quality of the generated CAD models, we first convert the B-Rep models into point clouds. Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models. We employ IoU, COV, MMD-CD, JSD, and NC as metrics to evaluate the quality of our generated CAD models [achlioptas2018learning]. IoU, MMD-CD, and NC have been introduced in Section 5.5.
@@ 6.2 Training Efficiency Comparison
Fusion 360 provides a server-side plugin called Fusion 360 Gym [willis2021fusion], which sequentially receives CAD modeling commands sent from external programs and returns the results to those programs via Hypertext Transfer Protocol (HTTP). Based on this plugin, we connect our reinforcement learning algorithm with Fusion 360 to enable closed-loop training.
We compared our novel training gym with Fusion 360 Gym over 10,000 training steps. Using parallel computation (1, 8, and 16 environments), we evaluated both memory consumption and execution time per step. Table 1 shows that Fusion 360 Gym consumes substantially more memory. The frequent use of the revert operation in Fusion 360 introduces a risk of memory overflow. In contrast, our gym employs a stable mark-and-revert mechanism that maintains steady memory usage without significant fluctuations. In addition, by replacing Fusion 360’s network communication mechanism with multi-process parallelism and shared memory architecture, we achieve a significant reduction in data transmission latency between the reinforcement learning algorithm and the gym. We get up to $39X$ speedup compared with Fusion 360 Gym, as shown in Table 2. More discussion of the limitations of Fusion 360 Gym can be found in  A.1.
@@
Table 2: Comparison of execution time per step (unit: ms). It shows that our method gets up to $39X$ speedup compared with Fusion 360 Gym.
@@ 6.3 B-Rep Modeling Comparison
    Graph Transformer Network (GTN) [shi2020masked]: Utilize multi-head self-attention with edge-aware positional encoding, effectively capturing long-range dependencies and global topological constraints.
@@ 6.5 Reward Ablation Study
Table 5: Reward function ablation study. It shows that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, while the metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
@@
Leveraging the ablated dataset curated in Section 6.3, we perform ablation studies to systematically analyze the impact of different reward function combinations on experimental results. Table 5 shows the outcomes of these studies. Reward functions such as IoU, MMD, and NC are used to measure geometric similarity. NR is used to capture the cosine similarity of high-dimensional B-Rep features. The experimental results show that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, meanwhile, the metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
@@ 6.6 Result Comparison of Reconstruction
Table 6: Quantitative results on the B-Rep based reconstruction task. Our method significantly outperforms Fusion 360 Gallery across all metrics on our test dataset (1kk models).
@@
Table 7: Quantitative comparison with cadrille and CAD-Recode on the reconstruction task, evaluated on the complete CAD-Recode validation dataset (1kk models). Our method significantly outperforms both baselines across all metrics.
@@
    Fusion 360 Gallery [willis2021fusion] extracts continuous geometric parameters from B-Rep models and uses imitation learning to generate command sequences, followed by a confidence-based local search for optimization.
@@
    CAD-Recode [rukhovich2024cad] translates a point cloud into Python code by leveraging a relatively small language model as a decoder and combined with a lightweight point cloud projector. The method is trained on a self-constructed, million-scale training dataset composed exclusively of sketch-extrusion operations.
@@
    cadrille [kolodiazhnyi2025cadrille] proposes a multi-modal CAD reconstruction model that can simultaneously process three input modalities, including point clouds, images, and text. The method adopts a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning fine-tuning using online feedback.
@@
Currently, most supervised learning methods are trained on datasets that include only sketch-extrusion operations. To overcome this limitation, we extend the output modules of Fusion 360 Gallery models to support revolve operations; more details can be found in  A.2. And we retrain the model on our dataset. As illustrated in Figure 6, our approach outperforms Fusion 360 Gallery in completeness and detail preservation. Quantitative results in Table 6 demonstrate substantial improvements across all evaluation metrics.
6.7 Comparison with Traditional Feature Recognition Methods
Figure 8: Comparison results with SolidWorks and Autodesk Inventor. The regions highlighted in magenta indicate features that cannot be recognized. We consider four representative cases.
To compare with traditional feature-recognition methods, we select the industrial CAD software Dassault SolidWorks [solidworks] and Autodesk Inventor [inventor]. Both provide state-of-the-art built-in feature-recognition plugins and can therefore be regarded as representative implementations of advanced feature-recognition techniques. Our experiments show that our method still exhibits several advantages over them, as shown in Figure 8:
(a) After two extrusion operations followed by a Boolean union, traditional feature-recognition methods fail to correctly interpret the new topology and instead recognize a single feature that does not correspond to the original modeling operations. In contrast, our method successfully recovers the two separate extrusion operations.
(b) For extrusion operations whose directions are not perpendicular to the sketch plane, both SolidWorks and Autodesk Inventor fail to recognize the feature (the red region in the figure), whereas our method can handle such cases correctly.
(c) For irregular extrusions and revolutions, traditional feature recognition typically relies on an analytic assumption that the end face can be traced back to a “planar end face + single sketch” configuration. This assumption easily breaks down when the end face is defined as up-to-surface/body/face with an offset, or when the side wall includes variable draft angles, leading to recognition failures. By contrast, our method can still correctly reconstruct the operations even with irregular profiles and non-uniform side walls.
(d) When multiple extrusion or revolution operations are combined using Boolean union/difference/intersection, the resulting model often exhibits complex topology, making it difficult for feature-recognition methods to recover the original “seed” extrusions/revolutions together with their sketches and directions; as a result, the editable feature history is lost. In comparison, our method can still decompose the merged result and reconstruct a plausible and consistent sequence of modeling operations.
@@ 6.8 Failure Case
We categorize the cases that our method currently cannot handle into three types. The first type arises from a complex model with many faces. Figure 9a exhibits the generation result of a complex gear model, in which some details are missing. Figure 10 depicts the quantitative relationship between generated model complexity (measured by face count) and various evaluation metrics. For simple models with a low face count, IoU and NC remain relatively high, but gradually decline as complexity increases. COV performs well on models with a low number of faces but declines significantly once the face count exceeds approximately 35–40. JSD, which is initially low, rises rapidly as the face count increases, especially beyond 40 faces. Overall, while the global reconstruction quality of our model remains acceptable beyond 35–40 faces, the fidelity of fine-grained geometric details deteriorates noticeably.
@@ 7.2 Future Work
Through extensive experiments, it has been found that our method still has shortcomings in the detailed feature learning of complex models and the supported operation amount. In the next phase, we will focus on developing detail-preserving feature extraction networks that support a broader range of input modalities and expanding the gym to support more commonly used modeling operations.
Appendix A Appendix
A.1 The Difficulty of Fusion 360 Framework Integration in Training Gym
[…]
A.2 Our Extension to Fusion 360 Gallery
We extend the method described in the Fusion 360 Gallery paper to support revolve operations, which requires two components: (i) extending the training network so that it can output revolve operations, and (ii) extending the Fusion 360 plugin to execute revolve commands. This plugin runs inside Fusion 360 and is responsible for executing the modeling commands sent from the client. Both components are required for Fusion 360 to fully support revolve operations in our framework.
More specifically, on the model-output side, we expand the action space from extrude-only to a unified mixed-action mode that supports both extrusion and revolution, while introducing only minimal changes to the inference interface and execution backend. The network now predicts a discrete action type $a\in{\text{extrude},\text{revolve}}$, a face selection, and a Boolean operation. For extrusion, the prediction remains a pair of parallel, non-coplanar planar faces $(f_{s},f_{e})$ together with a Boolean operation (newbody/intersection/union/subtraction). For revolution, the output is simplified to a single revolve-eligible face $f_{r}$ (non-planar, typically cylindrical, conical, spherical, or toroidal) plus a Boolean operation; the revolve axis is not regressed by the network but is instead inferred by the plugin described below.
The Fusion 360 plugin provides extended support for the revolve operation in Fusion 360. Given a selected surface of revolution, the system automatically infers a suitable rotation axis and generates a corresponding 2D sketch aligned with that axis. For different types of surfaces of revolution, the plugin adopts specialized construction logic:
  • 1.
    Cylindrical and conical surfaces: Feature points are extracted from the boundary in parametric space, projected onto the rotation axis, and then connected in sequence to form the sketch profile.
  • 2.
    Spherical surfaces: Boundary points are obtained from the parametric domain and connected to the sphere center to construct the revolve sketch.
  • 3.
    Toroidal surfaces: If the boundary in parametric space is a complete circle, this circle is directly used as the sketch. Otherwise, boundary points are connected to the center of the minor circle to build the sketch profile.
For surface types where a stable rotation axis or sketch cannot be inferred under the above rules, the plugin deliberately rejects processing them to avoid generating unreliable or meaningless features. All generated sketches and features are created within a dedicated reconstruction component to maintain structural consistency. If no valid entity can be found for reconstruction, the corresponding Boolean operations are safely skipped so as not to interfere with the overall modeling workflow.